Circular No. 6094

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVAE 1994ab, 1994ac T. Matheson and A. V. Filippenko, University of California at Berkeley,

report that preliminary inspection of an uncalibrated CCD spectrogram (range 420–700 nm, resolution 0.7 nm), obtained on Oct. 11 with the 3-m Shane reflector at Lick Observatory, suggests that SN 1994ab (IAUC 6089) is of type Ia, roughly 1–2 months past maximum, with a recession velocity of 10500 km/s. SN 1994ac (ibid.) is of type II, dominated by Hα emission, with a recession velocity of 5300 km/s.

GRS 1915+105 AND X-RAY NOVA IN SCORPIUS F. Nagase, H. Inoue, T. Kotani, and Y. Ueda, on behalf of the ASCA

Team, communicate: “We observed the two x-ray novae GRS 1915+105 and GRO J1655–40 with ASCA on Sept. 27–28, with exposure times of 20000 s each. Both sources were bright in the ASCA soft x-ray band during the observations. Preliminary analyses reveal that the superluminal x-ray transient source GRS 1915+105, which was observed during Sept. 27.03– 27.54 UT, showed an intensity variation that was small within ±10 percent of the average flux over time scales from minutes to hours. The average flux in the range 2–10 keV was 1.0 × 10^{−8} erg cm^{−2} s^{−1} (or ∼ 500 mCrab). The GIS and SIS spectra in the range 1–10 keV can be roughly represented by a power-law model with a photon index of ∼ 2.2 and an absorption column density of 4.3 × 10^{22} cm^{−2}, with significant spectral softening at high energy. A weak edge-like structure is visible around 7 keV; however, no conspicuous lines are seen in the spectra.

The x-ray nova in Scorpius (GRO J1655–40) was again observed with ASCA during Sept. 27.54–28.18. The intensity variation was small during the observation (< ±20 percent), in contrast with the large intensity variation in the previous observation on Aug. 23 (IAUC 6063). The average flux in the range 2–10 keV was 7.7 × 10^{−9} erg cm^{−2} s^{−1} (or ∼ 400 mCrab). The GIS and SIS spectra in the range 1–10 keV can be roughly represented by a power-law spectrum with a photon index of ∼ 2.0 and an interstellar absorption column density of 4.4 × 10^{21} cm^{−2}, with significant spectral softening at high energy. A weak but significant structure is seen around 6–7 keV.”

1994 October 12 Daniel W. E. Green